# On algebraic properties of some non-prime ideals of collections of cells

### v2

@@ 4. Cohen-Macaulay property and Hilbert series for closed path polyominoes having zig-zag walks
Observe that the result in Theorem 4.13 holds also for all collection of cells $\mathcal{Q}_{i}$ and $\mathcal{R}_{i,k}$ described in Set-up 4.2 and Set-up 4.8. In fact, if $\mathcal{R}$ is one of such configurations, by [3, Lemma 3.3] and by construction, we can argue that there exists a cell $A$ in $\mathcal{R}$ such that $\mathcal{R}\setminus\{A\}$ is simple. So, by [8, Proposition 4.9] (and [37, Corollary 3.1.7]) we have $\dim(K[\mathcal{R}])=|V(\mathcal{R})|-|\mathcal{R}|$. Therefore, we can obtain this claim using the same arguments developed in this section.
@@
Let $\mathcal{P}$ a closed path with a zig-zag walk. Then the $h$-polynomial of $K[\mathcal{P}]$ is the rook polynomial of $\mathcal{P}$ and $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$.
@@
Let $s_{1}=\max J_{\mathcal{Q}_{r}}$ and $s_{2}=\min J_{\mathcal{Q}_{r}}$. Following the collections of cells involved, observe that, $\mathcal{R}_{n_{s},s}$ is a zig-zag collection and and $\mathcal{S}_{1,s_{1}}=\mathcal{Q}_{r}$. In particular, by Corollary 3.8, $h_{K[\mathcal{R}_{n_{s},s}]}(t)$ is the rook polynomial of $\mathcal{R}_{n_{s},s}$. Therefore, if $k\in J_{\mathcal{Q}_{r}}$ and $i\in[n_{k}]$, it suffices to prove the following claim: if $h_{K[\mathcal{R}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{R}_{i,k}$ then $h_{K[\mathcal{S}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{S}_{i,k}$.
So, assume $h_{K[\mathcal{R}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{R}_{i,k}$. For all non-negative integer $k$, denote by $r^{(1)}_{k}$ and $r^{(2)}_{k}$, respectively, the coefficient of $t^{k}$ in the polynomials $h_{K[\mathcal{R}_{i,k}]}(t)$ and $h_{K[\mathcal{R}_{i,k}^{\prime}]}(t)$. Denote by $r_{k}=r^{(1)}_{k}+r^{(2)}_{k-1}$ (with the convention $r_{-1}^{(2)}=0$). Observe that $r_{k}$ is the coefficient of $t^{k}$ in $h_{K[\mathcal{S}_{i,k}]}(t)$. Hence, it suffices to show that $r_{k}$ is the number of $k$-rook configurations of $\mathcal{S}_{i,k}$. Recall that, as explained in the proof of Lemma 4.12, we can refer to Figure 20 and express $\mathcal{R}_{i,k}=\mathcal{S}_{i,k}\setminus\{A_{1}\}$ and $\mathcal{R}_{i,k}^{\prime}=\mathcal{S}_{i,k}\setminus(\{B,A_{1},\ldots,A_{r}\})$. So, consider that we can write $r_{k}=r_{1}+r_{2}$, where $r_{1}$ is the number of $k$-rook configurations of $\mathcal{S}_{i,k}$ having no rook placed in $A_{1}$ and $r_{2}$ is the number of $k$-rook configurations of $\mathcal{S}_{i,k}$ with a rook placed in $A_{1}$. It is not difficult to argue that $r_{1}$ is also equal to the number of $k$-rook configurations of $\mathcal{R}_{i,k}$, while $r_{2}$ is equal the number of $(k-1)$-rook configurations of $\mathcal{R}_{i,k}^{\prime}$. Therefore, $r_{1}=r_{k}^{(1)}$ and $r_{2}=r_{k-1}^{(2)}$, obtaining that $r_{k}$ is the number of $k$-rook configurations of $\mathcal{S}_{i,k}$. As consequence, $h_{K[\mathcal{R}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{S}_{i,k}$, proving our claim.
@@
Following the collections of cells involved and observing that $\mathcal{Q}_{0}=\mathcal{P}$, we can obtain our result if for all $i\in[r]$ we prove the following claim: if $h_{K[\mathcal{Q}_{i}]}(t)$ is the rook polynomial of $\mathcal{Q}_{i}$ then $h_{K[\mathcal{Q}_{i-1}]}(t)$ is the rook polynomial of $\mathcal{Q}_{i-1}$. But it is not difficult to see that this fact can be proved by the same argument used before, so we can conclude. ∎
@@
    From the previous cases it follows that every maximal interval of $\mathcal{P}$ has rank less or equal to $3$ and $\mathcal{P}$ does not contain any sub-polyomino as in Figure 14b. Hence the only possibility is that $\mathcal{P}$ is as in Figure 21. In such a case, Macaulay2 computations show that the reduced Hilbert series of $\mathcal{P}$ is given by
    $\frac{1+16\,t+100\,t^{2}+308\,t^{3}+486\,t^{4}+376\,t^{5}+134\,t^{6}+20\,t^{7}+t^{8}}{\left(1-t\right)^{16}}$
    so $h_{7}=20\neq 16=h_{1}$, which is a contradiction.
    Figure 21. A closed path.
@@
    Let
    $\mathrm{HP}_{K[\mathcal{P}]}(t)=\frac{h(t)}{(1-t)^{|V(\mathcal{P})|-|\mathcal{P}|}}$
    be the reduced Hilbert series of $K[\mathcal{P}]$. Then $h(t)$ is the rook polynomial of $K[\mathcal{P}]$ and $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$.
@@
We need to discuss just two cases. In the first one assume that $\mathcal{P}$ has a zig-zag walk. Then (1) and (2) follow from Theorems 4.13 and 4.15. Concerning (3), we know that $K[\mathcal{P}]$ is not Gorenstein from Proposition 4.17. Moreover, by [3, Proposition 6.1] $\mathcal{P}$ does not contain any $L$-configuration, so any two maximal inner intervals of $\mathcal{P}$ intersect themselves in the cells displayed in Figure 14a or 14b (up to reflections or rotations). Hence, we argue that if $\mathcal{P}$ has a zig-zag walk then it contains maximal blocks of length two. This means that (3) holds if $\mathcal{P}$ has zig-zag walks. Now, in the second case suppose that $\mathcal{P}$ does not contain a zig-zag walk. The claims (1) and (2) are already proved in [7, Theorem 5.5] and (3) in [7, Theorem 5.7]. In conclusion, the theorem is completely proved. ∎
5. Some open questions
We conclude this paper with some considerations and insights for future research. About the Cohen-Macaulayness of the inner $2$-minor ideals of collections of cells, it is known that if $\mathcal{P}$ is a simple collection of cells then $K[\mathcal{P}]$ is Cohen-Macaulay (see Proposition 1.3). Other classes of non-simple polyominoes whose coordinate ring is Cohen-Macaulay are discussed in [6], [21] and [26], although in these cases the examined classes consist just of prime polyominoes. In that cases, it is sufficient to study the Gröbner basis of the polyomino ideal by two well-known theorems of Sturmfels ([36]) and of Hochster ([2]). One of the main goal of this work is to investigate some classes of non-simple collections of cells whose coordinate ring is not a domain but it is Cohen-Macaulay. In particular, we proved that the zig-zag collections and the closed paths with zig-zag walks are Cohen-Macaulay, as well as the collections of cells $\mathcal{Q}_{i}$ and $\mathcal{R}_{i,k}$ introduced respectively in Set-up 4.2 and Set-up 4.8. Also, the coordinate ring of the collection of cells in Example 3.7 is Cohen-Macaulay. Nowadays, an example of a collection of cells having a not Cohen-Macaulay coordinate ring is still unknown. Therefore, the following general question naturally arises.
Question 5.1.
Let $\mathcal{P}$ be a collection of cells. Then $K[\mathcal{P}]$ is Cohen-Macaulay?
If $\mathcal{P}$ is a zig-zag collection we proved in Corollary 3.9(1) a necessary condition for $K[\mathcal{P}]$ to be Gorenstein. Motivated by some examples tested in Macaulay2, we ask if that condition is also sufficient, as mentioned in Remark 3.10. So, we formulate the following question:
Question 5.2.
Let $\mathcal{P}$ be a zig-zag collection supported by the intervals $I_{1},\dots,I_{\ell}$. Let $\mathcal{P}_{i}$ be the collection of cells related to $I_{i}$ and suppose that $\mathcal{P}_{i}$ is a square for all $i\in[\ell]$. Then, is $K[\mathcal{P}]$ Gorenstein?
Finally, in Theorem 4.18 we provide for the class of closed path polyominoes a description of the main algebraic properties of the related coordinate ring, characterizing them combinatorially. These properties are: be a domain, Hilbert-Poincaré series, regularity, be Cohen-Macaulay and Gorenstein. Recently in [32], the properties of Levelness and Pseudo-Gorenstein have been investigated in polyomino ideals for the first time, in particular for the class of path polyominoes (that are simple polyominoes). So another possible question is the following:
Question 5.3.
Let $\mathcal{P}$ be a closed path polyomino. Is it possible to characterize, by the combinatorics of $\mathcal{P}$, when $K[\mathcal{P}]$ is Level and Pseudo-Gorenstein?

### v3

@@ 4. Cohen-Macaulay property and Hilbert-Poincaré series of closed paths with zig-zag walks
Observe that the result in Theorem 4.13 holds also for all collection of cells $\mathcal{Q}_{i}$ and $\mathcal{R}_{i,k}$ described in Set-up 4.2 and Set-up 4.8. In fact, if $\mathcal{R}$ is one of such configurations, by [4, Lemma 3.3] and by construction, we can argue that there exists a cell $A$ in $\mathcal{R}$ such that $\mathcal{R}\setminus\{A\}$ is simple. So, by [9, Proposition 4.9] (and [37, Corollary 3.1.7]) we have $\dim(K[\mathcal{R}])=|V(\mathcal{R})|-|\mathcal{R}|$. Therefore, we can obtain this claim using the same arguments developed in this section.
@@
Let $\mathcal{P}$ be a closed path with a zig-zag walk. Then the $h$-polynomial of $K[\mathcal{P}]$ is the rook polynomial of $\mathcal{P}$ and $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$.
@@
Let $s_{1}=\max J_{\mathcal{Q}_{r}}$ and $s_{2}=\min J_{\mathcal{Q}_{r}}$. Following the collections of cells involved, observe that $\mathcal{R}_{n_{s},s}$ is a zig-zag collection and and $\mathcal{S}_{1,s_{1}}=\mathcal{Q}_{r}$. In particular, by Corollary 3.8, $h_{K[\mathcal{R}_{n_{s},s}]}(t)$ is the rook polynomial of $\mathcal{R}_{n_{s},s}$. Therefore, if $k\in J_{\mathcal{Q}_{r}}$ and $i\in[n_{k}]$, it suffices to prove the following claim: if $h_{K[\mathcal{R}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{R}_{i,k}$ then $h_{K[\mathcal{S}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{S}_{i,k}$.
So, assume $h_{K[\mathcal{R}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{R}_{i,k}$. For all non-negative integer $k$, denote by $r^{(1)}_{k}$ and $r^{(2)}_{k}$, respectively, the coefficient of $t^{k}$ in the polynomials $h_{K[\mathcal{R}_{i,k}]}(t)$ and $h_{K[\mathcal{R}_{i,k}^{\prime}]}(t)$. Denote by $r_{k}=r^{(1)}_{k}+r^{(2)}_{k-1}$ (with the convention $r_{-1}^{(2)}=0$). Observe that $r_{k}$ is the coefficient of $t^{k}$ in $h_{K[\mathcal{S}_{i,k}]}(t)$. Hence, it suffices to show that $r_{k}$ is the number of $k$-rook configurations of $\mathcal{S}_{i,k}$. Recall that, as explained in the proof of Lemma 4.12, we can refer to Figure 16 and express $\mathcal{R}_{i,k}=\mathcal{S}_{i,k}\setminus\{A_{1}\}$ and $\mathcal{R}_{i,k}^{\prime}=\mathcal{S}_{i,k}\setminus(\{B,A_{1},\ldots,A_{r}\})$. So, consider that we can write $r_{k}=r_{1}+r_{2}$, where $r_{1}$ is the number of $k$-rook configurations of $\mathcal{S}_{i,k}$ having no rook placed in $A_{1}$ and $r_{2}$ is the number of $k$-rook configurations of $\mathcal{S}_{i,k}$ with a rook placed in $A_{1}$. It is not difficult to argue that $r_{1}$ is also equal to the number of $k$-rook configurations of $\mathcal{R}_{i,k}$, while $r_{2}$ is equal the number of $(k-1)$-rook configurations of $\mathcal{R}_{i,k}^{\prime}$. Therefore, $r_{1}=r_{k}^{(1)}$ and $r_{2}=r_{k-1}^{(2)}$, obtaining that $r_{k}$ is the number of $k$-rook configurations of $\mathcal{S}_{i,k}$. As a consequence, $h_{K[\mathcal{R}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{S}_{i,k}$, proving our claim.
@@
Following the collections of cells involved and observing that $\mathcal{Q}_{0}=\mathcal{P}$, we can obtain our result if for all $i\in[r]$ we prove the following claim: if $h_{K[\mathcal{Q}_{i}]}(t)$ is the rook polynomial of $\mathcal{Q}_{i}$ then $h_{K[\mathcal{Q}_{i-1}]}(t)$ is the rook polynomial of $\mathcal{Q}_{i-1}$. However, it is not difficult to see that this fact can be proved by the same argument used before, so we can conclude. ∎
@@
    From the previous cases it follows that every maximal interval of $\mathcal{P}$ has a rank less than or equal to $3$ and $\mathcal{P}$ does not contain any sub-polyomino as in Figure 10b. Hence the only possibility is that $\mathcal{P}$ is as in Figure 3c. In such a case, Macaulay2 ([5, 13]) computations show that $h_{K[\mathcal{P}]}(t)=1+16\,t+100\,t^{2}+308\,t^{3}+486\,t^{4}+376\,t^{5}+134\,t^{6}+20\,t^{7}+t^{8},$ so $h_{7}=20\neq 16=h_{1}$, which is a contradiction.
@@
    The $h$-polyonomial of $K[\mathcal{P}]$ is the rook polynomial of $\mathcal{P}$ and $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$.
@@
We need to discuss just two cases. In the first one assume that $\mathcal{P}$ has a zig-zag walk. Then (1) and (2) follow from Theorems 4.13 and 4.15. Concerning (3), we know that $K[\mathcal{P}]$ is not Gorenstein from Proposition 4.17. Moreover, by [4, Proposition 6.1] $\mathcal{P}$ does not contain any $L$-configuration, so any two maximal inner intervals of $\mathcal{P}$ intersect themselves in the cells displayed in Figure 10a or 10b (up to reflections or rotations). Hence, we argue that if $\mathcal{P}$ has a zig-zag walk then it contains maximal blocks of length two. This means that (3) holds if $\mathcal{P}$ has zig-zag walks. Now, in the second case suppose that $\mathcal{P}$ does not contain a zig-zag walk. The claims (1) and (2) are already proved in [8, Theorem 5.5] and (3) in [8, Theorem 5.7]. In conclusion, the theorem is completely proved. ∎
We proved that the zig-zag collections and the closed paths with zig-zag walks are Cohen-Macaulay, as well as the collections of cells $\mathcal{Q}_{i}$ and $\mathcal{R}_{i,k}$ introduced respectively in Set-up 4.2 and Set-up 4.8. Nowadays, an example of a collection of cells having a not Cohen-Macaulay coordinate ring is still unknown. Therefore, from [19, 20, 31] and from the results of this work, the following general question naturally arises.
Question 4.19.
Let $\mathcal{P}$ be a collection of cells. Then, is $K[\mathcal{P}]$ Cohen-Macaulay?
Acknowledgement. The third author is supported by Scientific and Technological Research Council of Turkey TÜBİTAK under the Grant No: 122F128, and he is thankful to TÜBİTAK for their support. The first and the third author acknowledge support of INDAM-GNSAGA. The second author is supported by Scientific and Technological Research Council of Turkey TÜBİTAK under the Grant No: 118F169, and he is thankful to TÜBİTAK for their support. The second author also acknowledges the support of Erasmus funding during his stay at the University of Messina. Special thanks to Professor Rosanna Utano for her support and hospitality during this period.